OpenStudy (anonymous):

The product of the page numbers on two facing pages of a book is 156. Find the page numbers Medals will be given.

OpenStudy (saifoo.khan):

12 & 13

OpenStudy (anonymous):

thank you!

OpenStudy (saifoo.khan):

uw .: )

OpenStudy (anonymous):

im opening up a new assignment and then I will post more problems

OpenStudy (anonymous):

Your 1000th medal saifoo

OpenStudy (anonymous):

howdja do it saifooooooooo

OpenStudy (saifoo.khan):

Oh Yeah Imran...

OpenStudy (saifoo.khan):

DIdnt noticed! :D

OpenStudy (anonymous):

I gave you 1000th

OpenStudy (saifoo.khan):

Lol, Magic @paul

OpenStudy (saifoo.khan):

Oh, Cool, thnks for that. :)

OpenStudy (saifoo.khan):

i gave u your 1015th medal

OpenStudy (mathteacher1729):

But how did you arrive at your answer saifoo? "give someone a fish, feed them for a day... teach them to fish... feed them for life." :-p

OpenStudy (anonymous):

If you're looking for work shown: x(x+1)=156 x^2+x=156 x^2+x-156=0 (x+13)+(x-12)=0 (x+13)=0 or (x-12)=0 x=-13 or x=12 Pages can't be negative - therefore it's 12 and 12+1=13

OpenStudy (saifoo.khan):

Here u go @math

OpenStudy (anonymous):

oh, ok so magic then :)

OpenStudy (mathteacher1729):

Lord Bahel, great answer. :D +1 medal

OpenStudy (anonymous):

<3

OpenStudy (anonymous):

\[\{color}{green}{\text{really saifoo.khan}}\] \[\color{red}{\text{congratulations on 1000!}}\]

OpenStudy (anonymous):

Yes...spam every topic. Legit.

OpenStudy (anonymous):

\[\color{green}{\text{really saifoo.khan}}\] \[\color{red}{\text{congratulations on 1000!}}\] why can't i make this work!!

OpenStudy (anonymous):

Because, this site has pelletty coding. Why can't it use html5, phpbb codes, or anything the like?

OpenStudy (saifoo.khan):

THNKS! :D